OpenStudy (turingtest):

Find parametric equations for the tangent line at the point (cos(-4pi/6), sin(-4pi/6), -4pi/6) on the curve x = const , y = sin(t), z = t

OpenStudy (turingtest):

I got x=-1/2 y=sqrt(3)/2+t(cost) ????????? z=t-2pi/3

OpenStudy (jamesj):

Ok, so the displacement vector as a function of time is s(t) = (-1/2, sin t, t) and thus the velocity vector which is everywhere tangent is d/dt s(t) = v(t) = (0, cos t, 1) At the point P = (cos(-4pi/6), sin(-4pi/6), -4pi/6) , t = -4pi/6 = -2pi/3 Hence v(t) = v(-2pi/3) = (0, sqrt(3)/2, 1), call this D for direction.

OpenStudy (jamesj):

The line which is thus tangent at P has parametric form P + aD, for real numbers a

OpenStudy (jamesj):

or lambda, or whatever you like.

OpenStudy (turingtest):

I got x=-1/2 y=-sqrt(3)/2+a(cost) ????????? z=a-2pi/3

OpenStudy (jamesj):

Nooo. Note that we have calculated D explicitly for the t which corresponds to the point P.

OpenStudy (jamesj):

There should be absolutely no t in the expression. D = (0, sqrt(3)/2, 1)

OpenStudy (jamesj):

Hence the line P + aD = ....what?

OpenStudy (jamesj):

P = s(-4pi/6), that's where we got t from.

OpenStudy (jamesj):

and we use that value of t to evaluate the tangent direction v(t)

OpenStudy (jamesj):

D = v(-4pi/6)

OpenStudy (turingtest):

P=(-1/2,-sqrt(3)/2,-2pi/3) aD=(0,a*sqrt(3)/2,a) and then just add them?

OpenStudy (jamesj):

Yes. The parametric form of a line is a point on the line, P, plus a scalar times the direction of the line. E.g., (1,a,0) is a straight lin in the xy-plane that cross the x-axis at (1,0,0) and parallel to the y-axis because the direction of the line is (0,1,0)

OpenStudy (turingtest):

Ok, got it now, thanks a bunch. I need to take the Multivariable Calculus class on OCW apparently...

OpenStudy (jamesj):

good.

OpenStudy (anonymous):

how do you plug this in for each individual variable?

OpenStudy (anonymous):

such that x = y = z = t-4pi/6

OpenStudy (turingtest):

P=(-1/2,-sqrt(3)/2,-2pi/3) aD=(0,a*sqrt(3)/2,a) and then just add them: \[x=-1/2\]\[y=a \sqrt{3}/2-\sqrt{3}/2\]\[z=a-2 \pi/3\]

OpenStudy (turingtest):

Follow James' explanation to understand. But you said x=-1/2 was wrong, which James doesn't seem to agree with, so...?

OpenStudy (jamesj):

So D = v(t = -2pi/3) = (0, sin(t), 1) = (0, -sqrt(3)/2, 1)

OpenStudy (jamesj):

And P = (cos(-4pi/6), sin(-4pi/6), -4pi/6) = (-1/2, -sqrt(3)/2, -2pi/3)

OpenStudy (turingtest):

I thought I corrected that part already, but sonofa said the x=-1/2 part was wrong.

OpenStudy (jamesj):

hence the line has parametric form \[(-1/2, -\sqrt{3}/2(a + 1), a - 2\pi/3)\] Check every step, but I think this is right; I'm not doing this on paper and the probability of errors is higher. The x-ordinate is right because cos(-4pi/6) = -1/2 and the curve doesn't vary for x otherwise.

OpenStudy (jamesj):

I'd bring 1/6 outside the who thing to make it neater.

OpenStudy (turingtest):

P=(-1/2,-sqrt(3)/2,-2pi/3) aD=(0,-a*sqrt(3)/2,a) P+aD: x=−1/2 y=-a√3/2−√3/2=-sqrt(3)/2(a+1) z=a−2π/3 right, that's what I meant. I see the typo on the negative for the y, but you agree that x=-1/2, which was what sonofa said was wrong. He must have typed in x=1/2 or something. I don't know where he went though.

OpenStudy (anonymous):

i have instant check thing...i can check my answers immediately...i've tried 1/2 and -1/2 and both get marked incorrect

OpenStudy (anonymous):

says y is incorrect also

OpenStudy (jamesj):

Are you sure x = constant, and not x = cos(t)?

OpenStudy (turingtest):

James is an ex-professor of mathematics and came to the same conclusion as me, as can be seen above, so something is amiss here. Neither of us seem positive about the y part, but you can see we are both very confident that x=-1/2. I have my doubts about your program I guess.

OpenStudy (turingtest):

right, that's what I was thinking too

OpenStudy (turingtest):

I asked him at the beginning if he meant constant and he said yes

OpenStudy (anonymous):

the curve x=cost, y=sint, z=t

OpenStudy (anonymous):

that's copied and pasted

OpenStudy (anonymous):

it's not my program...it's what we do our homework on

OpenStudy (turingtest):

I asked you at the beginning if x=constant and you said yes!

OpenStudy (anonymous):

it's called webwork...you submit your answers and tells you if your right or wrong

OpenStudy (jamesj):

ah ha! s(t) = (cos t, sin t, t) So v(t) = (-sin t, cos t, 1) For t = -4pi/6 = -2pi/3, v(t) = (-sqrt(3)/2, -1/2, 1) = D Now P = (cos(-4pi/6), sin(-4pi/6), -4pi/6)) = (-1/2, -sqrt(3)/2,-2pi/3) Thus P + aD = (-1/2, -sqrt(3)/2,-2pi/3) + a(-sqrt(3)/2, -1/2, 1)

OpenStudy (anonymous):

wait...what's the x ?

OpenStudy (anonymous):

and what's the y?

OpenStudy (turingtest):

-1/2-a sqrt(3)/2

OpenStudy (jamesj):

Seriously Sonofa, can't you see how to pull out the x-ordinate here, or y-ordinate?

OpenStudy (anonymous):

i can but it looks like your still saying that x is -1/2 and i'm sayin it's getting marked incorrect when i submit that

OpenStudy (turingtest):

P+aD gotta add them dude!

OpenStudy (jamesj):

No, you're not reading correctly, which makes me very nervous about any verification you're doing.

OpenStudy (jamesj):

and what does the verification program expect for the parameter: a, t, lambda, .... ?

OpenStudy (anonymous):

t

OpenStudy (anonymous):

it's not a verification again....it's what our homework is submitted on

OpenStudy (turingtest):

whatever, then x=-1/2-t sqrt(3)/2

OpenStudy (jamesj):

P + tD = (-1/2, -sqrt(3)/2,-2pi/3) + t(-sqrt(3)/2, -1/2, 1) = (-1/2 - sqrt(3)/2 . t, -sqrt(3)/2 - t/2, - 2pi/3 + t)

OpenStudy (anonymous):

okay...now y and z are correct....x is still saying it's wrong

OpenStudy (jamesj):

Oh, yes, my fault -sin(-4pi/6) = sqrt(3)/2. Hence P + tD = (-1/2 + sqrt(3)/2 . t, -sqrt(3)/2 - t/2, - 2pi/3 + t)

OpenStudy (anonymous):

i got it!

OpenStudy (anonymous):

yes...exactly lol....it was + sqrt(3)/2 * t

OpenStudy (anonymous):

i could never thank you enough for helpign me with that

OpenStudy (anonymous):

would have never ever gotten that on my own

OpenStudy (anonymous):

both of you.....i owe you huge

OpenStudy (turingtest):

Nah, it's all thanks to James, as usual.